OpenStudy (anonymous):

solve. 1/(5x^2-5x) - 5/(5x-5)

OpenStudy (anonymous):

Equals what?

OpenStudy (anonymous):

oops, sorry =0

OpenStudy (anonymous):

x=0 i think...idk im tired so if i'm wrong sry

OpenStudy (anonymous):

\[\frac{1}{x(5x-5)}-\frac{5}{5x-5}=0\] \[\frac{1-5(x)}{x(5x-5)}=0\] \[1-5x=0\] \[5x=1\] \[x=\frac{1}{5}\]

OpenStudy (anonymous):

the answer would be 1-5x/5x(x-1), because in order to subtract you have to have the same denominator on both sides. So to make the denominator the same, you multiply the one on the right by x to make it 5x^2-5x which simplifies to 5x(x-1), just like the one on the left.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

wait... sorry haha did you want to solve for x?